# Birational complexity and conic fibrations

### v1

@@ 5.1. Complexity measures
where $r$ is a positive integer. In particular, $f_{i}(P)$ is a non-terminal valuation of $(X,B)$. By [28, Corollary 2.13], there exists a finite set $\mathcal{B}_{0}$ such that every non-terminal valuation $Q$ over $(X,B)$ satisfies $a_{Q}(X,B)\in\mathcal{B}_{0}$. Hence, $a_{f_{i}(P)}(X,B)\in\mathcal{B}_{0}$. Since $a_{P}(Y_{i},B_{i})<1$, by equality (5.2), we conclude that $a_{P}(Y_{i},B_{i})$ belongs to a finite set that only depends on $\mathcal{B}_{0}$. In particular, there is a lower bound $\epsilon_{0}>0$ for the coefficients of $B^{\prime}_{i}$ which is independent of $i$.
@@
If $f^{*}P$ is not a log canonical center of $(X,B)$, then we get a contradiction of the connectedness theorem [18, Theorem 1.2]. Hence, the divisor $f^{*}P$ appears with coefficient one in $B$. Since $Z$ is a Bott $\mathbb{Q}$-tower, we have that $\rho(Z)=\dim Z=n-1$ and $B_{Z}$ has $22n-$ 2 prime components. We conclude that $B$ has $2n$ prime components and $\rho(X)=\dim X=n$. Henceforth, the pair $(X,B)$ is toric by [11, Theorem 1.2].
@@ 5.2. Birational complexity and conic fibrations
[…]
@@
By condition (2), the variety $Y_{0,p}$ is a $\mathbb{Z}_{p}^{n-c}$-quasi-torsor over $X_{0}$ (see [9, Definition 2.7] and  [10, Proposition 4.10]). By condition (1), we know that $Y_{0,p}\rightarrow X_{0}$ only ramifies along $B_{0}$. By condition (3), if $p$ is a large prime number, then coefficients of each $\alpha_{i}D_{i}/p$ are fractional provided that $\vec{\alpha}\neq 0$ in $\mathbb{Z}_{p}^{n-c}$. In particular, the induced morphism $Y_{0,p}\rightarrow W$ has connected fibers for $p\gg 0$. Let $(Y_{0,p},B_{Y_{0,p}})$ be the log pull-back of $(X_{0},B_{0})$. By the above considerations, the pair $(Y_{0,p},B_{Y_{0,p}})$ is log Calabi–Yau. Since $Y_{0,p}\rightarrow X_{0}$ only ramifies along $\lfloor B_{0}\rfloor=B_{0}$, we conclude that $Y_{0,p}\rightarrow W$ is a Fano type morphism (see Lemma 2.17). By construction, the group $\mathbb{Z}_{p}^{n-c}$ acts on $(Y_{0,p},B_{Y_{0,p}})$ fiberwise over $W$. Thus, if $p\geq p_{0}(n)$, then we can apply Theorem 4.1. From now on, we just fix $p\gg 0$ and set $(Y_{0},B_{Y_{0}}):=(Y_{0,p},B_{Y_{0,p}})$. By Theorem 4.1, there exists a $\mathbb{Z}_{p}^{n-c}$-equivariant crepant birational model $(Y^{\prime},B_{Y^{\prime}})$ of $(Y_{0},B_{Y_{0}})$ over $W$ that admits a tower of Mori fiber spaces
@@
[…]
@@
We conclude that the birational complexity of $X_{d}$ is at most $d-1$. Then, the statement follows from Theorem 1.1. ∎
@@ 5.3. Dimension of the dual complex
[…]
@@ 5.4. Fundamental groups and alteration conic complexity
By the proof of [38, Theorem 3.3], we know that there is an isomorphism
@@
where $A_{\vec{m}}$ is a finite abelian group of rank at most $n-k_{\rm alt}(X,B)$ and $|N_{\vec{m}}|\leq c(n)$. Here, $c(n)$ is a constant that only depends on the dimension $n$. Write $G_{\vec{m}}:=\pi_{1}^{\rm reg}(X,\Delta_{\vec{m}})$. Let $Y_{\vec{m}}\rightarrow X$ be the cover associated to $G_{\vec{m}}$. Let $(Y_{\vec{m}},D_{\vec{m}})$ be the log pull-back of $(X,\Delta_{\vec{m}})$. By [8, Theorem 3], there is a normal finite abelian subgroup $A_{\vec{m}}\leqslant G_{\vec{m}}$ of index at most $c_{0}(n)$. By Lemma 2.17, the variety $Y_{\vec{m}}$ is of Fano type. We know that $A_{\vec{m}}$ has rank at most $n$. Write $A_{\vec{m}}:=\bigoplus_{i=1}^{n}\mathbb{Z}_{b_{i,\vec{m}}}$ with $b_{1,\vec{m}}\mid\dots\mid b_{n,\vec{m}}$. Let $s\in\{1,\dots,n\}$ be the smallest positive for which $b_{s,\vec{m}}\geq m_{0}(n,\{0,1\})$, where $m_{0}(n,\{0,1\})$ is as in Theorem 4.1. We replace $A_{\vec{m}}:=\bigoplus_{i=1}^{n}\mathbb{Z}_{b_{i,\vec{m}}}$ with $\bigoplus_{i=s}^{n}\mathbb{Z}_{b_{i,\vec{m}}}$. Note that the former group has index bounded above by $m_{0}(n,\{0,1\})^{s-1}$ in $A_{\vec{m}}$. After such replacement, by Theorem 4.1, we have an $A_{\vec{m}}$-equivariant crepant birational model $(Y^{\prime}_{\vec{m}},\Delta^{\prime}_{\vec{m}})$ of $(Y_{\vec{m}},\Delta_{\vec{m}})$ and a sequence of equivariant Mori fiber spaces:
[…]
We write $\psi_{1,j}:=\psi_{j}\circ\dots\circ\psi_{1}$. By construction, we know that there is a subset $I\subseteq\{1,\dots,r+1\}$ with $|I|=n-s$ such that each $\psi_{i}$ with $i\in I$ is a conic fibration. Let $A_{\vec{m},i}$ be the homomorphic image of $A_{\vec{m}}$ acting on $Z_{i}$. Furthermore, for each $i\in I$, the boundary divisor $\lfloor\Delta_{i}\rfloor$ has two disjoint $A_{\vec{m},i}$-invariant prime components that are horizontal over $Z_{i+1}$. This implies that $k(Y^{\prime}_{\vec{m}},\Delta^{\prime}_{\vec{m}})\leq s.$ Hence, the rank of $A_{\vec{m}}$ is at most $n-k(Y^{\prime}_{\vec{m}},\Delta^{\prime}_{\vec{m}})$. So the rank of $A_{\vec{m}}$ is at most $n-k_{\rm alt}(X,B)$. Therefore, we have a short exact sequence as in (5.7).
@@ 5.5. Log Calabi–Yau pairs of coregularity 0
Log Calabi–Yau pairs of coregularity zero are of special interest [16]. Indeed, every $n$-dimensional Fano variety of coregularity zero $X$ admits a $2$-complement $(X,B)$ of coregularity zero (see, e.g., [14, Theorem 4]). Thus, a $n$-dimensional Fano variety $X$ of coregularity zero admits a $1$-complement of coregularity zero $(X,B)$ after possibly replacing $X$ with a $2$-to-$1$ cover. By [31, Theorem 1.4], we know that $c_{\rm bir}(X,B)\leq n$. If $c_{\rm bir}(X,B)<n$, then Theorem 1.1 implies that $(X,B)$ admits a non-trivial birational Mori fiber space. Thus, the building blocks of log Calabi–Yau pairs of coregularity zero and index one are:
@@
[…]
@@
[…]
@@
[…]
@@
Assume that $\phi_{1}$ is a conic fibration over $X_{1}$ and $B^{\prime}$ has a prime component that maps $2$-to-$1$ onto $X_{1}$. Up to a $2$-to-$1$ cover of $X^{\prime}$, possibly ramified along $B^{\prime}$, we may assume that $B^{\prime}$ has two prime components that are sections over $X_{1}$. Hence, we are in the situation of the previous paragraph. In this case, we conclude that $c_{\rm alt}(X,B)=0$, so condition (a) holds.
@@
As mentioned above, we have $c_{\rm alt}(X,B)\in\{0,1,\dots,n\}$ for a $n$-dimensional log Calabi–Yau pair of index one and coregularity zero. Theorem 5.13 implies that the values $1$ and $2$ do not occur for $c_{\rm alt}(X,B)$ in dimension $2$. For $3$-folds, only the values 00 and $1$ are known to occur (see Question 6.8).

### v2

@@ 5. Proof of the theorems
5.1. Finite actions on rationally connected varieties
In this subsection, we prove a geometric Jordan property for finite actions on rationally connected varieties.
Proof of Theorem 1.1.
We proceed by induction on the dimension. Let $n$ be the dimension of $X$. The statement for $n=1$ by using the Jordan property for $\mathbb{P}{\rm GL}(2,\mathbb{K})$. By using the Jordan property [40], we may assume that $G$ is indeed abelian. Let $X\dashrightarrow X^{\prime}$ be a $G$-equivariant birational map with a $G$-equivariant Fano type morphism $X^{\prime}\rightarrow Z$ for which $\dim(X^{\prime}/Z)$ is minimized. Let $G_{f}<G$ be the normal subgroup acting fiberwise over $Z$. Note that $X^{\prime}$ admits an $G$-invariant NN complement $(X^{\prime},B^{\prime})$ over $Z$, so we are in the situation of Theorem 4.1. Let $m_{0}(n,\frac{1}{N}\mathbb{Z}\cap[0,1])$ be as in the statement of Theorem 4.1. If $|G_{f}|\leq m_{0}(n,\frac{1}{N}\mathbb{Z}\cap[0,1])$, then we may replace $G$ with $G^{|G_{f}|}$ and the statement follows by induction on the dimension. If $|G_{f}|>m_{0}(n,\frac{1}{N}\mathbb{Z}\cap[0,1])$, then we can apply Theorem 4.1 with $W=Z$  [^4] The proof applies as long as we have a relative complement over $W$.. Let $G_{0}\leqslant G$ be the normal subgroup of bounded index given by Theorem 4.1. Here, the index is bounded in terms of $n$ and $N$. There are two possibilities:
  1. (i)
    If $\psi_{0}$ is a strict conic fibration, then $X^{\prime}\simeq_{\rm bir}\mathbb{P}^{1}\times X_{0}$ is an $G_{0}$-equivariant isomorphism. In this case the statement follows by induction on the dimension.
  2. (ii)
    If $\psi_{0}$ is not a strict conic fibration.
If (ii) holds, then we replace $G$ with $G_{0}$ so $\dim(X^{\prime}/Z)$ drops. We can iterate the argument and (ii) can only hold finitely many times as in each iteration we replace $A$ with a subgroup of bounded index and the dimension of $X^{\prime}/Z$ drops. Thus, eventually, we are in the situation of (i) so the statement follows from induction on the dimension. ∎
@@ 5.2. Complexity measures
where $r$ is a positive integer. In particular, $f_{i}(P)$ is a non-terminal valuation of $(X,B)$. By [26, Corollary 2.13], there exists a finite set $\mathcal{B}_{0}$ such that every non-terminal valuation $Q$ over $(X,B)$ satisfies $a_{Q}(X,B)\in\mathcal{B}_{0}$. Hence, $a_{f_{i}(P)}(X,B)\in\mathcal{B}_{0}$. Since $a_{P}(Y_{i},B_{i})<1$, by equality (5.2), we conclude that $a_{P}(Y_{i},B_{i})$ belongs to a finite set that only depends on $\mathcal{B}_{0}$. In particular, there is a lower bound $\epsilon_{0}>0$ for the coefficients of $B^{\prime}_{i}$ which is independent of $i$.
@@
If $f^{*}P$ is not a log canonical center of $(X,B)$, then we get a contradiction of the connectedness theorem [18, Theorem 1.2]. Hence, the divisor $f^{*}P$ appears with coefficient one in $B$. Since $Z$ is a Bott $\mathbb{Q}$-tower, we have that $\rho(Z)=\dim Z=n-1$ and $B_{Z}$ has 2n22n-2 prime components. We conclude that $B$ has $2n$ prime components and $\rho(X)=\dim X=n$. Henceforth, the pair $(X,B)$ is toric by [11, Theorem 1.2].
@@ 5.3. Birational complexity and conic fibrations
    ${X^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{0}}$${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{1}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{r-1}}$${X_{r}}$
@@
By condition (2), the variety $Y_{0,p}$ is a $\mathbb{Z}_{p}^{n-c}$-quasi-torsor over $X_{0}$ (see [8, Definition 2.7] and  [9, Proposition 4.10]). By condition (1), we know that $Y_{0,p}\rightarrow X_{0}$ only ramifies along $B_{0}$. By condition (3), if $p$ is a large prime number, then coefficients of each $\alpha_{i}D_{i}/p$ are fractional provided that $\vec{\alpha}\neq 0$ in $\mathbb{Z}_{p}^{n-c}$. In particular, the induced morphism $Y_{0,p}\rightarrow W$ has connected fibers for $p\gg 0$. Let $(Y_{0,p},B_{Y_{0,p}})$ be the log pull-back of $(X_{0},B_{0})$. By the above considerations, the pair $(Y_{0,p},B_{Y_{0,p}})$ is log Calabi–Yau. Since $Y_{0,p}\rightarrow X_{0}$ only ramifies along $\lfloor B_{0}\rfloor=B_{0}$, we conclude that $Y_{0,p}\rightarrow W$ is a Fano type morphism (see Lemma 2.17). By construction, the group $\mathbb{Z}_{p}^{n-c}$ acts on $(Y_{0,p},B_{Y_{0,p}})$ fiberwise over $W$. Thus, if $p\geq p_{0}(n)$, then we can apply Theorem 4.1. From now on, we just fix $p\gg 0$ and set $(Y_{0},B_{Y_{0}}):=(Y_{0,p},B_{Y_{0,p}})$. By Theorem 4.1, there exists a $\mathbb{Z}_{p}^{n-c}$-equivariant crepant birational model $(Y^{\prime},B_{Y^{\prime}})$ of $(Y_{0},B_{Y_{0}})$ over $W$ that admits a tower of Mori fiber spaces
@@
${(Y_{0},B_{Y_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{/\mathbb{Z}_{p}^{n-c}}$${(Y^{\prime},B_{Y^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{/\mathbb{Z}_{p}^{n-c}}$$\scriptstyle{\phi_{0}}$${Y_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$$\scriptstyle{/\mathbb{Z}_{p}^{k_{1}}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r-1}}$${Y_{r}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{/\mathbb{Z}_{p}^{k_{r}}}$$\scriptstyle{\phi_{r}}$${(X,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X_{0},B_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X^{\prime},B^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{0}}$${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{1}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{r-1}}$${X_{r}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{r}}$${W}$
@@
We conclude that the birational complexity of $X_{d}$ is at most d1d-1. Then, the statement follows from Theorem 1.2. ∎
@@ 5.4. Dimension of the dual complex
(5.4) ${Y^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${Y_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{2}}$${Y_{2}}$
@@ 5.5. Fundamental groups and alteration conic complexity
By the proof of [33, Theorem 3.3], we know that there is an isomorphism
@@
where $A_{\vec{m}}$ is a finite abelian group of rank at most $n-k_{\rm alt}(X,B)$ and $|N_{\vec{m}}|\leq c(n)$. Here, $c(n)$ is a constant that only depends on the dimension $n$. Write $G_{\vec{m}}:=\pi_{1}^{\rm reg}(X,\Delta_{\vec{m}})$. Let $Y_{\vec{m}}\rightarrow X$ be the cover associated to $G_{\vec{m}}$. Let $(Y_{\vec{m}},D_{\vec{m}})$ be the log pull-back of $(X,\Delta_{\vec{m}})$. By [7, Theorem 3], there is a normal finite abelian subgroup $A_{\vec{m}}\leqslant G_{\vec{m}}$ of index at most $c_{0}(n)$. By Lemma 2.17, the variety $Y_{\vec{m}}$ is of Fano type. We know that $A_{\vec{m}}$ has rank at most $n$. Write $A_{\vec{m}}:=\bigoplus_{i=1}^{n}\mathbb{Z}_{b_{i,\vec{m}}}$ with $b_{1,\vec{m}}\mid\dots\mid b_{n,\vec{m}}$. Let $s\in\{1,\dots,n\}$ be the smallest positive for which $b_{s,\vec{m}}\geq m_{0}(n,\{0,1\})$, where $m_{0}(n,\{0,1\})$ is as in Theorem 4.1. We replace $A_{\vec{m}}:=\bigoplus_{i=1}^{n}\mathbb{Z}_{b_{i,\vec{m}}}$ with $\bigoplus_{i=s}^{n}\mathbb{Z}_{b_{i,\vec{m}}}$. Note that the former group has index bounded above by $m_{0}(n,\{0,1\})^{s-1}$ in $A_{\vec{m}}$. After such replacement, by Theorem 4.1, we have an $A_{\vec{m}}$-equivariant crepant birational model $(Y^{\prime}_{\vec{m}},\Delta^{\prime}_{\vec{m}})$ of $(Y_{\vec{m}},\Delta_{\vec{m}})$ and a sequence of equivariant Mori fiber spaces:
${(Y^{\prime}_{\vec{m}},\Delta^{\prime}_{\vec{m}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{1}}$${(Z_{1},\Delta_{1},\mathbf{M}_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{2}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{r}}$${(Z_{r},\Delta_{r},\mathbf{M}_{r})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{r+1}}$${{\rm Spec}(\mathbb{K})}$
We write $\psi_{1,j}:=\psi_{j}\circ\dots\circ\psi_{1}$. By construction, we know that there is a subset $I\subseteq\{1,\dots,r+1\}$ with |I|=ns|I|=n-s such that each $\psi_{i}$ with $i\in I$ is a conic fibration. Let $A_{\vec{m},i}$ be the homomorphic image of $A_{\vec{m}}$ acting on $Z_{i}$. Furthermore, for each $i\in I$, the boundary divisor $\lfloor\Delta_{i}\rfloor$ has two disjoint $A_{\vec{m},i}$-invariant prime components that are horizontal over $Z_{i+1}$. This implies that $k(Y^{\prime}_{\vec{m}},\Delta^{\prime}_{\vec{m}})\leq s.$ Hence, the rank of $A_{\vec{m}}$ is at most $n-k(Y^{\prime}_{\vec{m}},\Delta^{\prime}_{\vec{m}})$. So the rank of $A_{\vec{m}}$ is at most $n-k_{\rm alt}(X,B)$. Therefore, we have a short exact sequence as in (5.7).
@@ 5.6. Log Calabi–Yau pairs of coregularity 0
Log Calabi–Yau pairs of coregularity zero are of special interest [16]. Indeed, every $n$-dimensional Fano variety of coregularity zero $X$ admits a $2$-complement $(X,B)$ of coregularity zero (see, e.g., [14, Theorem 4]). Thus, a $n$-dimensional Fano variety $X$ of coregularity zero admits a $1$-complement of coregularity zero $(X,B)$ after possibly replacing $X$ with a 22-to-11 cover. By [31, Theorem 1.4], we know that $c_{\rm bir}(X,B)\leq n$. If $c_{\rm bir}(X,B)<n$, then Theorem 1.2 implies that $(X,B)$ admits a non-trivial birational Mori fiber space. Thus, the building blocks of log Calabi–Yau pairs of coregularity zero and index one are:
@@
${X_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{2}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r}}$${X_{r}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r+1}}$${{\rm Spec}(\mathbb{K}).}$
@@
${(Y_{0},\Delta_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{1}}$$\scriptstyle{\pi_{1}}$${(Y_{1},\Delta_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{2}}$${Y_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{3}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{s}}$${Y_{s}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X_{0},B_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${X_{1}}$
@@
${X^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{2}}$${X_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{3}}$${{\rm Spec}(\mathbb{K}),}$
@@
Assume that $\phi_{1}$ is a conic fibration over $X_{1}$ and $B^{\prime}$ has a prime component that maps 22-to-11 onto $X_{1}$. Up to a 22-to-11 cover of $X^{\prime}$, possibly ramified along $B^{\prime}$, we may assume that $B^{\prime}$ has two prime components that are sections over $X_{1}$. Hence, we are in the situation of the previous paragraph. In this case, we conclude that $c_{\rm alt}(X,B)=0$, so condition (a) holds.
@@
As mentioned above, we have $c_{\rm alt}(X,B)\in\{0,1,\dots,n\}$ for a $n$-dimensional log Calabi–Yau pair of index one and coregularity zero. Theorem 5.13 implies that the values $1$ and $2$ do not occur for $c_{\rm alt}(X,B)$ in dimension $2$. For 33-folds, only the values 0 and $1$ are known to occur (see Question 6.8).